# Task Arithmetic with Support Languages for Low-Resource ASR

### v1

@@ Abstract
The development of resource-constrained approaches to automatic speech recognition (ASR) is of great interest due to its broad applicability to many low-resource languages for which there is scant usable data. Existing approaches to many low-resource natural language processing tasks leverage additional data from higher-resource languages that are closely related to a target low-resource language. One increasingly popular approach uses task arithmetic to combine models trained on different tasks to create a model for a task where there is little to no training data. In this paper, we consider training on a particular language to be a task, and we generate task vectors by fine-tuning variants of the Whisper ASR system. For pairings of high- and low-resource languages, we merge task vectors via a linear combination, optimizing the weights of the linear combination on the downstream word error rate on the low-resource target language’s validation set. We find that this approach consistently improves performance on the target languages. [^1] Code for this project can be found at https://github.com/ddegenaro/mozilla-asr-challenge.
@@ 1 Introduction
In this paper, we explore the use of task arithmetic (Nagasawa et al., 2025) to leverage higher-resource languages for low-resource ASR. We fully fine-tune the pre-trained whisper-tiny model Radford et al. (2023) and train Low-Rank Adapters (LoRA) Hu et al. (2022) for whisper-large-v3 for each target language. To improve performance in low-resource settings, we incorporate information from genetically related higher-resource “support languages.” For each target language, we identify closely related languages and train corresponding models on Common Voice (CV) data Ardila et al. (2020) using the same training setup. We then apply task arithmetic to combine target and support models (or adapters), selecting merged models based on Word Error Rate (WER) on the target development sets. Our results show consistent improvements across languages, demonstrating the effectiveness of support language integration for low-resource ASR with task arithmetic.
@@ 2.2 Low-Resource ASR
A complementary line of work uses task vectors (TVs) (Ilharco et al., 2023), defined as parameter differences between fine-tuned and base models, which can be composed to transfer capabilities across tasks or domains. This “task arithmetic” can also be achieved via combining LoRA layers (Zhang et al., 2023; Chitale et al., 2023; Chronopoulou et al., 2024; Li et al., 2025). Recent studies apply task arithmetic to ASR, demonstrating that combining task vectors from related higher-resource languages with target-language vectors improves performance under limited data (Ramesh et al., 2024; Su et al., 2024; Nagasawa et al., 2025).
@@ 3.1 Models and Training
We use LoRA layers to fine-tune whisper-large-v3, and we apply full fine-tuning to whisper-tiny to fit within the constraints of Task 2, which calls for a model that is less than 500MB in size. Due to resource constraints, we quantize whisper-large-v3 to 4-bit. We employ early stopping based on the WER on the validation set. The hyperparameters used for training can be seen in Table 1 in Appendix A.
@@ 3.2 Leveraging Related Languages
We make use of supplemental data from various languages which are genetically related to the languages being considered in this shared task. Following Nagasawa et al. 2025, we train task vectors (or LoRA adapter-based equivalents) for each support language and then combine them with the target language models at evaluation time.
Let $\theta$ be the weights of the pretrained Whisper model and $\theta_{S}$ be the weights of a model fine-tuned on ASR data from the support language S{S}. We define the TV $\tau_{S}$ as:
$\tau_{S}=\theta_{S}-\theta$ (1)
Let $\theta_{T}$ be the weights of the model trained on ASR data in the target language. We then apply $\tau_{S}$ to $\theta_{T}$ to create the final model $\theta_{final}$:
$\theta_{final}=\theta_{T}+\lambda\tau_{S}$ (2)
@@
Support languages were chosen based on genetic relatedness to the target language, shared scripts, and the availability of scripted speech data in Common Voice. We fine-tune whisper-tiny and train LoRA adapters for whisper-large-v3 separately for each target language and each support language grouping. Whisper’s language detection is disabled and the language token is fixed to a Whisper proxy language: a pre-training language selected to maximize token overlap with the target language while sharing the same script. This initialization aims to start adaptation from a representation space as close as possible to the target language. Mappings between target languages, support languages, proxy languages, and scripts are provided in Table 2, with language codes listed in Tables 3 and 4. The selected “high-resource” familial languages were often still relatively low-resource. In such cases, we combine multiple related languages when possible. For example, for Puno Quechua, we use 13 Quechuan languages to construct the support models. [^2] We were unable to identify suitable familial languages for Papantla Totonac and Toba Qom, and therefore do not use any support languages for these targets. For Scots, the closest familial language is English; in this case, we likewise omit support languages and rely on the fact that Whisper’s pre-training is heavily dominated by English.
After training all models, we tune the task-arithmetic scaling parameter $\lambda$ on the validation set for each target language. Prior work has shown that performance is sensitive to the choice of $\lambda$ Li et al. (2025); Nagasawa et al. (2025). Instead of a grid search, we use Bayesian optimization with ten evaluations over the range $[0,1]$. Preliminary experiments indicated that this range was unnecessarily large, as all optimal values of $\lambda$ were below $0.5$.
@@ 3.3 Data Sources and Munging
We use standard scripted speech data from CV for all support languages as well as for the five unseen target languages (Adyghe, Basaa, Kabardian, Puno Quechua, and Ushojo). For the remaining 21 target languages, we use the task-provided the spontaneous speech data.
@@ 4 Results
Results for full fine-tuning whisper-tiny on the development split can be see in Figure 1. The development data for the unseen languages comes from the CV scripted speech corpus. We see improvements from incorporating the support language across all of the target languages, improving upon the target model by up to -0.092 with an average WER improvement of -0.0367. WER can be understood as the percentage of incorrect words in the transcript. Thus, depending on the performance of the original model, a 3-9% decrease in WER can indicate a large improvement in an ASR model. Compared to the task baseline results on the development split, our small model improves upon 4 of the languages: Ushojo, Toba Qom, Basaa, and Kabardian. We expect the WER to be slightly higher on the test split, specifically for the unseen language data which will be in a different domain (spontaneous speech instead of scripted speech).
@@
Results from the whisper-large model can be found in Figure 2. For the many of languages, this model performed worse than both the task baseline and whisper-tiny, only improving upon the development baseline for Ushojo. We only submitted whisper-large transcripts for Scots, Papantla Totonac, and Betwi to task 1, as those were the only models that outperformed whisper-tiny. Additionally, there were instances in which the support language was not helpful for the target language whisper-large model. This indicates that some of these adapters were not trained well and did not appear to capture much information from the target and support languages. Despite this, the incorporation of the support language led to average WER improvements of -0.054, and in the case of Adyghe this method improved upon the base model by -0.497. This indicates that task arithmetic can be successfully performed via linear combinations of LoRA layers and is not limited to TVs.
Figure 2: Incorporating the support language into the whisper-large via linear combinations of LoRA layers led to WER improvements, though overall performance is worse than whisper-tiny.
@@ 5 Discussion
[…]
Finally, the task baseline based on fine-tuning the Massively Multilingual Speech (MMS) model Pratap et al. (2024) outperforms most of our systems. This suggests that MMS may be a more suitable base model for low-resource ASR. Additionally, it is likely that some of our choices for support languages were suboptimal, and the quantitative method for determining language similarity described above might have helped us to find better support languages.
@@ 6 Conclusion
We describe the system for our submission to the 2025 Mozilla Common Voice Spontaneous Speech challenge. We fine-tune whisper-tiny and train LoRA adapters for whisper-large on each target language and use task arithmetic with genetically related support languages sharing a script with target languages to improve upon each base model. We find that this approach consistently lowers the WER on the target languages, particularly when the target language model performs poorly alone. Rather than using the largest model possible, our results show that a more targeted approach using smaller models and relevant support data can lead to greater success for low-resource ASR.
@@
We were unable to complete a search for the optimal training hyperparameters due to lack of a compute resources and time. In addition to training hyperparameters, the search space for $\lambda$ was relatively broad: $[0,1]$. In the end, all of the optimal values were <0.4<0.4. Even within the search space of [0,4], any small change in $\lambda$ varied the performance significantly. For future work, this search space should be narrowed in order to allow for the optimal $\lambda$ to be discovered. Furthermore, as discussed in Section 5, the choice of base model significantly impacted the overall results of this experiment. For the best performance, a better base model should be chosen and a full hyperparameter sweep should be run for language specific training values as well as $\lambda$ values.
@@ Appendix A Hyperparameter Settings
Batch size whisper-large 32
@@
Appendix B Languages
bxk Bantu, Zone J lg Swahili Latin
cgg Bantu, Zone J lg Swahili Latin
koo Bantu, Zone J lg Swahili Latin
lke Bantu, Zone J lg Swahili Latin
ruc Bantu, Zone J lg Swahili Latin
ttj Bantu, Zone J lg Swahili Latin
rwm Bantu, Zone D lg Swahili Latin
bas* Bantu, Zone A lg Swahili Latin
kcn Afro-Asiatic, Semitic mt Maltese Latin
led Nilo-Saharan, E. Sudanic, Central luo, kln Somali Latin
lth Nilo-Saharan, E. Sudanic, W. Nilotic luo, kln Somali Latin
ukv Nilo-Saharan, E. Sudanic, E. Nilotic luo, kln Somali Latin
hch Uto-Aztecan ncx, nhi, nlv, yaq, tar, var Spanish Latin
meh Oto-Manguean, Mixtecan cut, cux, mau Spanish Latin
mmc Oto-Manguean, Otomian cut, cux, mau Spanish Latin
top Totonacan $\emptyset$ Spanish Latin
qxp* Quechuan qu{p,x,r,y,s}, qx{a,u,w,t}, qv{l,a}, qw{a,s} Spanish Latin
tob Guaicuruan $\emptyset$ Spanish Latin
aln Indo-European, Albanian sq Albanian Latin
el-CY Indo-European, Greek el Greek Greek
sco Indo-European, Germanic, West $\emptyset$ English Latin
ush* Indo-European, Indo-Aryan, Dardic ur Urdu Perso-Arabic
ady* Northwest Caucasian ab Kazakh Cyrillic & Latin
kbd* Northwest Caucasian ab Kazakh Cyrillic & Latin
bew Austronesian, Malay id, ms Malay Latin
pne Austronesian, Malayo-Polynesian, Kenyah id, ms Malay Latin
Table 2: Low-resource target languages, their selected support counterparts and Whisper proxy language codes. * indicates a “test-only” language.

### v2

@@ Abstract
The development of resource-constrained approaches to automatic speech recognition (ASR) is of great interest due to its broad applicability to many low-resource languages for which there is scant usable data. Existing approaches to many low-resource natural language processing tasks leverage additional data from higher-resource languages that are closely related to a target low-resource language. One increasingly popular approach uses task arithmetic to combine models trained on different tasks to create a model for a task where there is little to no training data. In this paper, we consider training on a particular language to be a task, and we generate task vectors by fine-tuning variants of the Whisper ASR system. For pairs of high- and low-resource languages, we merge task vectors via a linear combination which is optimized on the downstream word error rate on the low-resource target language’s validation set. Across 23 low-resource target languages for which we evaluate this technique, we find consistent word error rate improvements of up to 10% compared to a baseline without our approach. [^1] Code for this project can be found at https://github.com/ddegenaro/mozilla-asr-challenge.
@@ 1 Introduction
In this paper, we explore the use of task arithmetic (Ilharco et al., 2023) to leverage higher-resource languages for low-resource ASR. We fully fine-tune the pre-trained whisper-tiny model Radford et al. (2023) and train Low-Rank Adapters (LoRA) Hu et al. (2022) for whisper-large-v3 for each target language. To improve performance in low-resource settings, we incorporate information from genetically related higher-resource “support languages” Nagasawa et al. (2025). For each target language, we identify closely related languages and train corresponding models on Common Voice (CV) data Ardila et al. (2020) using the same training setup. We then apply task arithmetic to combine target and support models (or adapters), selecting merged models based on WER on the target development sets. Our results show consistent improvements across languages, demonstrating the effectiveness of support language integration for low-resource ASR with task arithmetic.
@@ 2.2 Low-Resource ASR
A complementary line of work uses task vectors (TVs; Ilharco et al., 2023), defined as the parameter-wise difference between fine-tuned and base models, which can be composed to transfer capabilities across tasks or domains. This “task arithmetic” can also be achieved via combining LoRA layers (Zhang et al., 2023; Chitale et al., 2023; Chronopoulou et al., 2024; Li et al., 2025). Recent studies apply task arithmetic to ASR (Ramesh et al., 2024; Su et al., 2024), demonstrating its utility in resource-constrained domains. Nagasawa et al. (2025) also use task arithmetic to improve low-resource language ASR models using high-resource counterparts, but their experiment was limited to four target languages which had all been briefly included in Whisper’s training data. For this shared task, we expand their approach to more languages and language families, where most of the languages were not present during Whisper training.
@@ 3.1 Models and Training
We use LoRA layers to fine-tune whisper-large-v3, and we apply full fine-tuning to whisper-tiny to fit within the constraints of Task 3. Due to resource constraints, we quantize whisper-large-v3 to 4-bits. We employ early stopping based on the WER on the validation set. The hyperparameters used for training can be seen in Table 1 in Appendix A.
@@ 3.2 Leveraging Related Languages
We make use of supplemental data from various languages which are genetically related to the languages being considered in this shared task. Following Ilharco et al. (2023) and Nagasawa et al. (2025), we train task vectors (or LoRA adapter-based equivalents) for each support language and then combine them with the target language models at evaluation time.
Let $\boldsymbol{\theta}$ be the weights of the pretrained Whisper model and $\boldsymbol{\theta}_{S}$ be the weights of a model fine-tuned on ASR data from the support language(s) S{S}. We define the TV $\boldsymbol{\tau}_{S}$ as:
$\boldsymbol{\tau}_{S}=\boldsymbol{\theta}_{S}-\boldsymbol{\theta}$ (1)
Note that $\boldsymbol{\theta}_{S}$ may be optimized jointly over several support languages taken in tandem. Let $\boldsymbol{\theta}_{T}$ be the weights of the model trained on ASR data in the target language. We then apply $\boldsymbol{\tau}_{S}$ to $\boldsymbol{\theta}_{T}$ to create the final model $\boldsymbol{\theta}_{final}$:
$\boldsymbol{\theta}_{final}=\boldsymbol{\theta}_{T}+\lambda\boldsymbol{\tau}_{S}$ (2)
@@
Support languages were chosen based on genetic relatedness to the target language, shared scripts, and the availability of scripted speech data in CV. We did not attempt any metric-based selection of support languages, as data availability in CV was already a significant limitation. We fine-tune whisper-tiny and train LoRA adapters for whisper-large-v3 separately for each target language and each support language grouping. Whisper’s language detection is disabled and the language token is fixed to a Whisper proxy language: a pre-training language selected to maximize token overlap with the target language while sharing the same script. This initialization aims to start adaptation from a representation space as close as possible to the target language. Mappings between target languages, support languages, proxy languages, and scripts are provided in Table 4, with language codes listed in Tables 2 and 3. The selected “high-resource” familial languages were often still relatively low-resource. In such cases, we combine multiple related languages when possible. For example, for Puno Quechua, we use 13 Quechuan languages to construct the support models. [^2] We were unable to identify suitable familial languages for Papantla Totonac and Toba Qom, and therefore do not use any support languages for these targets. For Scots, the closest familial language is English; in this case, we likewise omit support languages and rely on the fact that Whisper’s pre-training is heavily dominated by English.
After training all models, we tune the task-arithmetic scaling parameter $\lambda$ on the validation set for each target language. Prior work has shown that performance is sensitive to the choice of $\lambda$ Li et al. (2025); Nagasawa et al. (2025). Instead of a grid search, we use Bayesian optimization with ten evaluations over the range $[0,1]$.
@@ 3.3 Data Sources and Munging
We use standard scripted speech data from CV for all support languages and for the five unseen target languages (Adyghe, Basaa, Kabardian, Puno Quechua, and Ushojo). For the remaining 21 target languages, we use the task-provided spontaneous speech data.
@@ 4 Results
Results for full fine-tuning whisper-tiny on the development split can be see in Figure 1. The development data for the unseen languages comes from the CV scripted speech corpus. We see improvements from incorporating the support language across all of the target languages, improving upon the target model by up to -9.93% with an average WER improvement of -3.51%. WER can be understood as the percentage of incorrect words in the transcript. Thus, depending on the performance of the original model, up to a 3-9% decrease in WER can indicate a large improvement in an ASR model. While some language-specific improvements are small, the worst-case scenario of this approach is falling back on the baseline of only using the target language model. Experimenting with addition of a support language TV can only help. Compared to the task baseline results on the development split, our small model improves upon four of the languages: Ushojo, Toba Qom, Basaa, and Kabardian. We expect the WER to be slightly higher on the test split, specifically for the unseen language data which will be in a different domain (spontaneous speech instead of scripted speech).
@@
Results from the whisper-large-v3 model can be found in Figure 2. For the many of languages, this model performed worse than both the task development baseline and whisper-tiny. We only submitted whisper-large-v3 transcripts for Scots, Papantla Totonac, and Betwi to task 1, as those were the only models that outperformed whisper-tiny. Additionally, there were instances in which the support language was not helpful for the target language whisper-large-v3 model. This indicates that some of these adapters were not trained well and did not appear to capture much information from the target and support languages. Despite this, the incorporation of the support language led to average WER improvements of -5.43%, and in the case of Adyghe this method improved upon the base model by -49.72%. This indicates that task arithmetic can be successfully performed via linear combinations of LoRA layers and is not limited to TVs.
Figure 2: Incorporating the support language into the whisper-large-v3 via linear combinations of LoRA layers led to WER improvements, though overall performance is worse than whisper-tiny.
@@ 5 Discussion
Figure 3: There is a moderate correlation between the cosine similarity of each target language and its support language and the impact of the support language TV on the target model. The shaded area is the 95% CI.
@@
Finally, the task baseline based on fine-tuning the Massively Multilingual Speech (MMS) model Pratap et al. (2024) outperforms most of our systems. This suggests that MMS may be a more suitable base model than Whisper for low-resource ASR. Additionally, it is likely that some of our choices for support languages were suboptimal, and the quantitative method for determining language similarity described above might have helped us to find better support languages.
@@ 6 Conclusion
We describe the system for our submission to the 2025 Mozilla Common Voice Spontaneous Speech challenge. We fine-tune whisper-tiny and train LoRA adapters for whisper-large-v3 on each target language and use task arithmetic with genetically related support languages to improve upon each target model. We find that this approach consistently lowers the WER on the target languages, particularly when the target language model performs poorly alone. Rather than using the largest model possible, our results show that a more targeted approach using smaller models and relevant support data can lead to greater success for low-resource ASR.
@@
We were unable to complete a search for the optimal training hyperparameters due to lack of a compute resources and time. In addition to training hyperparameters, the search space for $\lambda$ was relatively broad: $[0,1]$. In the end, all of the optimal values were <0.5<0.5. Even within the search space of [0, 0.5], any small change in $\lambda$ varied the performance significantly. For future work, this search space should be narrowed in order to allow for the optimal $\lambda$ to be discovered. Furthermore, as discussed in Section 5, the choice of base model significantly impacted the overall results of this experiment. For the best performance, a better base model should be chosen and a full hyperparameter sweep should be run for language specific training values as well as $\lambda$ values.
@@
Finally, future work can improve selection criteria/prioritization of support languages and data, including comparisons of phonology, lexical borrowings, dictionary overlap, or phylogenetic tree distance.
Acknowledgments
We thank Joe Garman and Christopher Cervantes for their thoughtful feedback. These experiments were conducted using Georgetown HPC.
@@ Appendix A Hyperparameter Settings
See Table 1.
Appendix B Languages
Language codes can be found in Tables 2 and 3. Support languages for the low-resource languages are reported in Table 4.
@@ Appendix B Languages
Batch size whisper-large-v3 32
@@
bxk Bantu, Zone J lg Swahili Latin
cgg Bantu, Zone J lg Swahili Latin
koo Bantu, Zone J lg Swahili Latin
lke Bantu, Zone J lg Swahili Latin
ruc Bantu, Zone J lg Swahili Latin
ttj Bantu, Zone J lg Swahili Latin
rwm Bantu, Zone D lg Swahili Latin
bas* Bantu, Zone A lg Swahili Latin
kcn Afro-Asiatic, Semitic mt Maltese Latin
led Nilo-Saharan, E. Sudanic, Central luo, kln Somali Latin
lth Nilo-Saharan, E. Sudanic, W. Nilotic luo, kln Somali Latin
ukv Nilo-Saharan, E. Sudanic, E. Nilotic luo, kln Somali Latin
hch Uto-Aztecan ncx, nhi, nlv, yaq, tar, var Spanish Latin
meh Oto-Manguean, Mixtecan cut, cux, mau Spanish Latin
mmc Oto-Manguean, Otomian cut, cux, mau Spanish Latin
top Totonacan $\emptyset$ Spanish Latin
qxp* Quechuan qu{p,x,r,y,s}, qx{a,u,w,t}, qv{l,a}, qw{a,s} Spanish Latin
tob Guaicuruan $\emptyset$ Spanish Latin
aln Indo-European, Albanian sq Albanian Latin
el-CY Indo-European, Greek el Greek Greek
sco Indo-European, Germanic, West $\emptyset$ English Latin
ush* Indo-European, Indo-Aryan, Dardic ur Urdu Perso-Arabic
ady* Northwest Caucasian ab Kazakh Cyrillic & Latin
kbd* Northwest Caucasian ab Kazakh Cyrillic & Latin
bew Austronesian, Malay id, ms Malay Latin
pne Austronesian, Malayo-Polynesian, Kenyah id, ms Malay Latin
Table 4: Low-resource target languages, their selected support counterparts and Whisper proxy language codes. * indicates a “test-only” or “unseen” language. Data sourced from each language’s Wikipedia page.